DETECTION OF SiH_{4}–H_{2}O BY FOURIER TRANSFORM MICROWAVE SPECTROSCOPY

\underline{Y. KAWASHIMA}, Department of Applied Chemistry, Kanagawa Institute of Technology, Atsugi, Kanagawa 243-03, Japan; R. D. SUENRAM, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899; E. HIROTA, The Graduate University of Advanced Studies, Hayama, Kanagawa 240-01, Japan.

The microwave spectra of the normal species, ^{29}Si, ^{30}Si, ^{18}O, and ^{2}H SiH_{4}–H_{2}O have been measured using a pulsednozzle Fourier transform microwave (FTMW) spectrometer. Two, four and five strong transitions were observed for the J = 1 ← 0, 2 ← 1, and 3 ← 2 transitions, respectively in the 7 to 22 GHz region. Almost all of the observed lines consisted of two or three components. The spectral patterns observed for SiH_{4}–H_{2}O were different from those of CH_{4}–H_{2}O^{a}, CH_{4}– HCl^{b}, and Ar–SiH_{4}^{c}. The observed line frequencies were fitted to the expression for the rotational spectrum of a linear molecule. The centrifugal distortion constant of the set appearing at lower frequencies than the K=0 transition is much smaller than those of other transitions. The rotational constants thus obtained, give an Si-O bond length of 3.38Å, which is much shorter than the C-O bond length in CH_{4}–H_{2}O (3.70Å). The value of the stretching force constant estimated for SiH_{4}–H_{2}O from the rotational and centrifugal distortion constants is larger than that for the C-O bond in CH_{4}–H_{2}O. We thus conclude that the Si-O bond in SiH_{4}–H_{2}O is much stronger than the C-O bond in CH_{4}–H_{2}O. From the observed Stark effect in SiH_{4}–H_{2}O, the electric dipole moment was determined to be 1.730 Debye. The structural and dipole data indicates that the water moiety is located with its C_{2} axis coinciding with the van der Waals bond.

aR.D. Suenram, G.T. Fraser, F.J. Lovas, and Y. Kawashima, J. Chem. Phys., 101, 7230 (1994)

bY. Ohshima and Y. Endo, J. Chem. Phys., 93, 6256 (1990)

cY. Kawashima, R.D. Suenram, G.T. Fraser, F.J. Lovas, and Y. Hirota, 51^{st} Symposium on Molecular Spectroscopy, (1996).

TD08